OpenStudy (anonymous):

Simplify: [3(n+3)!]/5n!

OpenStudy (anonymous):

there are no parentheses around the number right? it is \[\frac{3(n+3)!}{5n!}\]

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

you distribute the 3 first I believe to get 3n +3!

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

oh hell no

OpenStudy (anonymous):

for example \[3\times 5!=3\times 5\times 4\times 3\times 2\times 1\]

OpenStudy (anonymous):

don't distribute anything leave the \(\frac{3}{5}\) out front

OpenStudy (anonymous):

then \[(n+3)!=(n+3)(n+2)(n+1)n!\] cancel the \(n!\) top and bottom and you are left with only \[(n+3)(n+2)(n+1)\]

OpenStudy (anonymous):

it takes a bit of getting used to we can do an example with numbers

OpenStudy (anonymous):

if you don't distribute the 3 why is it n+3 and not n+1?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

ok lets make sure it is clear what \((n+3)!\) means